A rectangular lawn 90 ft by 120 is mowed, cutting uniform (equal) widths around the outside. How wide a strip has been cut when the lawn is half mowed?
Answer by edjones(8007)

90*120=10800
10800/2=5400 area when 1/2 the lawn is mowed.
(90-x)(120-x)=5400
10800-210x+x^2=5400
x^2-210x+5400=0
(x-30)(x-180)=0
x=30 ft. strip has been cut when the lawn is half mowed.
.
Ed